# An Explicit M1 Radiation-hydrodynamics Scheme for Three-dimensional Protostellar Evolution

### v1

@@ 5.4 Step 4: radiative transfer step
where the superscripts $\mathrm{R}$ and $\mathrm{L}$ denote the values of physical quantities in the right and left cells adjacent to the interface, respectively. As explained in Section 2.2, we need to solve the RT of the RSLA component using only the streaming photons $E^{\mathrm{R}}_{\mathrm{rad,S}}$ to reproduce the physical diffusion by the numerical diffusion represented by the second term in Equation (82). Therefore, the variables $(E,F)$ in Equation (82) correspond to ($E^{\mathrm{R}}_{\mathrm{rad,S}},\bm{F}_{\mathrm{S}}$) and ($\bm{F}_{\mathrm{S}},\bm{\mathsf{P}}_{\mathrm{rad,S}}$) for solving Equations (80) and (81). For the flux of the non-RSLA component, $(c/3\hat{\chi}\rho)\nabla E^{\mathrm{N}}_{\mathrm{rad}}$ in Equation (80), we compute the intercell numerical flux with the central difference method as
@@
Here, we estimate the value of $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ by distributing $E^{\mathrm{con}}_{\mathrm{rad}}$ based on Equations (16) and (49). However, as mentioned in Section 4.2, to capture the physical behavior of radiation escaping from an optically thick cell into an optically thin one, we must ensure that the streaming photons $E^{\mathrm{R}}_{\mathrm{rad,S}}$ dominate even in an optically thick cell if its neighboring cell is optically thin. To achieve this, we obtain $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ using the smaller of the optical depths of the two adjacent cells as
@@
5.5 Step 5: radiative momentum exchange step
@@ 6 NUMERICAL TESTS
In this section, we present a series of numerical tests to validate the newly developed RHD scheme. In Section 6.1, we validate the non-RSLA component introduced in Section 4.1, with respect to energy conservation (Section 6.1.1), radiation transport (Sections 6.1.2 and 6.1.3), and radiation pressure (Section 6.1.4). In Section 6.2, we demonstrate that the method introduced in Section 4.2, which incorporates information about the optical depth of neighboring cells, enables accurate treatment of RT even in regions with steep optical depth gradients. In Section 6.3, we verify that our spectral reconstruction method, described in Section 4.3, can track the evolution of the radiation spectrum consistently with the emitting sources, including its spatial variation. In all of the following tests, we set the reduction factor of the RSLA $f_{c}$ to $10^{-3}$.
6.1 Tests of the non-RSLA component
6.1.1 Energy conservation test during gas-radiation interaction
@@
Figure 1: Energy conservation test during gas-radiation interaction. The left and right panels show the results without and with the inclusion of the non-RSLA radiation component $E^{\mathrm{N}}_{\mathrm{rad}}$, respectively. The upper panels display the time evolution of the gas temperature, and the lower panels show the evolution of the energy density of each component. In the upper panels, the gray dashed line indicates the analytic equilibrium temperature when the system evolves while conserving its total energy (see text). In the lower panels, the blue solid line represents the total energy (gas plus radiation), the orange and green dashed lines denote the gas and radiation energy densities, respectively, and the red dotted line indicates the radiation energy density corresponding to the Planck distribution, $aT^{4}$.
@@
As discussed in Section 3.1, the use of the RSLA can lead to violations of energy conservation during the energy exchange between gas and radiation, making it difficult to accurately follow the evolution of optically thick, radiation-dominated systems. To address this issue, in Section 4.1, we have introduced a non-RSLA radiation component, $E^{\mathrm{N}}_{\mathrm{rad}}$, which is not subject to the RSLA and becomes dominant in optically thick regions such that $\tau_{\mathrm{cell}}\gtrsim 2/(3f_{c})$. To verify the effectiveness of this approach, we evolve a single computational cell over time under local emission and absorption processes, without including hydrodynamics or RT. This test demonstrates that our new scheme conserves the total energy and accurately reproduces the thermal equilibrium state expected under optically thick conditions.
@@
6.1.2 Radiative transfer test with the diffusion approximation
Figure 2: Radiative transfer test with the diffusion approximation. The blue solid line represents the spherically averaged radial profile in the steady-state from our simulation. The orange dashed line denotes the analytic solution. The gray dotted line indicates the computational cell size, which corresponds to the finite volume over which radiation is injected.
@@ 6.1.2 Radiative transfer test with the diffusion approximation
When radiation from a central source propagates through a uniform medium, the steady-state solution can be obtained analytically. In the optically thick limit, radiation propagates with diffusion as $F=-c\nabla E_{\mathrm{rad}}/(3\rho\sigma^{\nu}_{0})$. On the other hand, assuming a steady state, the flux can be expressed with the luminosity of the central radiation source $L$ as $F=L/(4\pi r^{2})$. As mentioned in J. Rosdahl & R. Teyssier (2015), combining the two expressions for the flux yields the radiation energy density of
@@
6.1.3 Advecting radiation pulse test
Figure 3: Advecting radiation pulse test. The top and middle panels show the distributions of density and temperature. The blue solid and orange dashed lines represent the static and moving cases, respectively, while the gray dotted lines denote the initial conditions. The bottom panel presents the relative errors between the static and moving cases. The green and red lines indicate the relative errors in density and temperature, respectively. All quantities are evaluated at $t=4.8\times 10^{-4}$ s.
@@ 6.1.3 Advecting radiation pulse test
where $T_{0}=10^{7}$ K, $T_{1}=2\times 10^{7}$ K, $\rho_{0}=1.2~\mathrm{g~cm}^{-3}$, $w=24$ cm, and $\mu=2.23$. The resulting initial profiles are shown as gray dotted lines in the top and middle panels of Figure 3. As the system evolves, the central radiation peak gradually broadens due to diffusion. We perform this test under two conditions: one in which the gas remains static, and another in which the gas moves uniformly at a constant velocity of $10^{5}~\mathrm{cm~s^{-1}}$. If the scheme accurately captures radiation advection in the dynamic diffusion limit, both cases should exhibit identical evolution in the comoving frame (M. R. Krumholz et al., 2007b). We set the computational domain to be a cubic box with a side length of $1024$ cm, resolved by $256^{3}$ uniform cells. Periodic boundary conditions are applied in all directions. We adopt a constant absorption opacity of $\kappa^{\nu}_{0}=10^{4}~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and neglect scattering. Under these conditions, the optical depth per cell is $\tau_{\mathrm{cell}}\gtrsim 4\times 10^{3}$, indicating that the radiation energy is dominated by the non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$.
@@
6.1.4 Radiation pressure tube test
Figure 4: Radiation pressure tube test. In the top panel, the blue and orange lines show the density $\rho$ and temperature $T$ profiles along the $x$-axis after ten sound-crossing times, while the dashed and dotted lines represent the analytic hydrostatic equilibrium solutions for density and temperature, respectively. In the bottom panel, we show the relative errors of the numerical results presented in the top panel with respect to the analytic solutions. The blue and orange lines represent the relative errors in density and temperature, respectively.
Here, we demonstrate that our scheme accurately captures radiation pressure in the optically thick regime, where the non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$ dominates. To this end, we examine whether our scheme can maintain a hydrostatic structure supported by both gas and radiation pressure. First, to construct the initial condition for 3D simulations, we obtain the gas profile in steady-state hydrostatic equilibrium under optically thick conditions, where radiation is transported by diffusion, by solving the following 1D equations, which describe the so-called radiation pressure tube problem (M. R. Krumholz et al., 2007b):
@@ 6.1.4 Radiation pressure tube test
6.2 Radiation transfer test with steep optical depth gradient
Figure 5: Radiation transfer test with steep optical depth gradient. We show spherically averaged radial profiles of outward luminosity in the steady state. The blue line represents the result obtained by accounting for the cell optical depth $\tau_{\mathrm{cell}}$ of adjacent cells, while the orange line shows the result obtained using only the local value of $\tau_{\mathrm{cell}}$. The horizontal dashed line marks the expected luminosity, $4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$. The vertical dotted line indicates the clump radius, $R_{\mathrm{c}}$.
As mentioned in Section 4.2, if radiation is divided into the RSLA and non-RSLA components solely based on the local value of $\tau_{\mathrm{cell}}$, it cannot escape from optically thick cells even when the adjacent cell is optically thin. To address this issue, we distribute the radiation components by also considering the $\tau_{\mathrm{cell}}$ values of neighboring cells, as detailed in Section 5.3 and 5.4. In this section, we verify that this method allows radiation to escape from optically thick regions into adjacent optically thin regions. To isolate the radiative behavior, we disable hydrodynamics and neglect scattering for simplicity.
We set up a computational box with a side length of $200~\mathrm{au}$, resolved with $128^{3}$ uniform cells. At the center of the box, we place a gas clump with radius $R_{\mathrm{c}}=30~\mathrm{au}$, density $\rho_{\mathrm{c}}=10^{-2}~\mathrm{g}~\mathrm{cm}^{-3}$, and temperature $T_{\mathrm{c}}=10^{5}~\mathrm{K}$. Outside the clump, we place a gas with a density of $\rho=10^{-20}~\mathrm{g}~\mathrm{cm}^{-3}$ and an initial temperature of $T=100~\mathrm{K}$. The gas temperature within the clump is kept fixed throughout the calculation. We set a constant absorption opacity of $\kappa^{\nu}_{0}=10^{2}~\mathrm{cm^{2}~g^{-1}}$. Due to this opacity value, the optical depth per cell becomes extremely large inside the clump, reaching $\tau_{\mathrm{cell}}=5\times 10^{11}$, while the region outside the clump remains optically thin. We then follow the radiation emitted by the clump. In the ideal case, the clump radiates from its surface with a luminosity of $L=4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$.
@@ 6.2 Radiation transfer test with steep optical depth gradient
6.3 Radiation spectrum test
In our scheme, the radiation spectrum is reconstructed from the photon mean energy, as described in Section 4.3. In this section, we validate the ability of our method to accurately track the evolution of the radiation spectrum. As illustrated by the dashed line in Figure 6, we place two clumps along the $x$-axis at positions $x=80$ and 80-80 au, with temperatures of $T_{\mathrm{c}}=6\times 10^{4}$ and $3\times 10^{4}$ K, and radii of $R_{\mathrm{c}}=10$ and $40$ au, respectively. Both clumps have identical luminosities, given by $L=4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$. The gas density inside each clump is set to $10^{-2}~\mathrm{g~cm}^{-3}$, while the ambient medium has a density of $10^{-16}~\mathrm{g~cm}^{-3}$ and an initial temperature of $100$ K. To isolate radiation effects, we turn off hydrodynamics and fix the gas temperature within the clumps. The opacity is set such that $\kappa^{\nu}_{0}=100~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and $\sigma^{\nu}_{0}=0~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ inside the clumps, and $\kappa^{\nu}_{0}=0~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and $\sigma^{\nu}_{0}=100~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ outside. This setup ensures that radiation is emitted only within the clumps, while outside it is scattered without absorption or emission. Starting from an initial condition with no radiation field, we evolve the system until it reaches a steady state.

### v2

@@ V.4 Step 4: radiative transfer step
where the superscripts $\mathrm{R}$ and $\mathrm{L}$ denote the values of physical quantities in the right and left cells adjacent to the interface, respectively. As explained in Section II.2, we need to solve the RT of the RSLA component using only the streaming photons $E^{\mathrm{R}}_{\mathrm{rad,S}}$ to reproduce the physical diffusion by the numerical diffusion represented by the second term in Equation (83). Therefore, the variables $(E,F)$ in Equation (83) correspond to ($E^{\mathrm{R}}_{\mathrm{rad,S}},\bm{F}_{\mathrm{S}}$) and ($\bm{F}_{\mathrm{S}},\bm{\mathsf{P}}^{\mathrm{R}}_{\mathrm{rad,S}}$) for solving Equations (81) and (82). For the flux of the non-RSLA component, $(c/3\hat{\chi}\rho)\nabla E^{\mathrm{N}}_{\mathrm{rad}}$ in Equation (81), we compute the intercell numerical flux with the central difference method as
@@
Here, we estimate the value of $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ by distributing $E^{\mathrm{con}}_{\mathrm{rad}}$ based on Equations (16) and (49). However, as mentioned in Section IV.2, to capture the physical behavior of radiation escaping from an optically thick cell into an optically thin one, we must ensure that the streaming photons $E^{\mathrm{R}}_{\mathrm{rad,S}}$ dominate even in an optically thick cell if its neighboring cell is optically thin. To achieve this, we obtain $E^{\mathrm{R}}_{\mathrm{rad,S}}$ and $E^{\mathrm{N}}_{\mathrm{rad}}$ using the smaller of the optical depths of the two adjacent cells as
@@
V.5 Step 5: radiative momentum exchange step
@@ VI NUMERICAL TESTS
In this section, we present a series of numerical tests to validate the newly developed RHD scheme. In Section VI.1, we validate the non-RSLA component introduced in Section IV.1, with respect to energy conservation (Section VI.1.1), radiation transport (Sections VI.1.2 and VI.1.3), radiation pressure (Section VI.1.4). In Section VI.2, we demonstrate that the method introduced in Section IV.2, which incorporates information about the optical depth of neighboring cells, enables accurate treatment of RT even in regions with steep optical depth gradients. In Section VI.3, we verify that our spectral reconstruction method, described in Section IV.3, can track the evolution of the radiation spectrum consistently with the emitting sources, including its spatial variation. Moreover, in Section VI.4, we measure the computational time of each computational operation in the test described in Section VI.3 and summarize the overhead of our new scheme.
In all of the following tests, we set the reduction factor of the RSLA $f_{c}$ to $10^{-3}$ and CFL number to $0.5$ for both the hydrodynamics and the M1 CFL condition. We summarize the main parameters of each test in Table 4.
Table 4: Parameters for the numerical tests. We adopt $f_{c}=10^{-3}$ and a CFL number of 0.5 for all tests.
Section Box Size Resolution $\kappa_{0}^{\nu}$ $\sigma_{0}^{\nu}$ $\mathrm{BC}_{\mathrm{hydro}}$ $\mathrm{BC}_{\mathrm{rad}}$
($\mathrm{cm}^{2}~\mathrm{g}^{-1}$) ($\mathrm{cm}^{2}~\mathrm{g}^{-1}$)
VI.1.1 $10^{11}$ cm 1 10 0
VI.1.2 $200$ au $32^{3}$ 0 $4.3\times 10^{-2}$ optically thick
VI.1.3 $1024$ cm $256^{3}$ $10^{4}$ 0 peridodic periodic
[…]
VI.2 $200$ au $128^{3}$ $10^{2}$ 0 optically thin
[…]
Notes. $\mathrm{BC}_{\mathrm{hydro}}$ and $\mathrm{BC}_{\mathrm{rad}}$ denote the outer boundary conditions for hydrodynamics and radiation, respectively.
The term “optically thick” refers to a boundary condition in which the radiation energy density decreases as $r^{-1}$
and the flux as $r^{-2}$. The term “optically thin” refers to a boundary condition where both the energy density
and flux decrease as $r^{-2}$.
∗1 Boundary conditions for xx/($y$, $z$) directions are denoted.
∗2 Opacity values inside/outside clumps are denoted.
VI.1 Tests of the non-RSLA component
VI.1.1 Energy conservation test during gas-radiation interaction
@@
Figure 1: Energy conservation test during gas-radiation interaction. The left and right panels show the results without and with the inclusion of the non-RSLA radiation component $E^{\mathrm{N}}_{\mathrm{rad}}$, respectively. The upper panels display the time evolution of the gas temperature, and the lower panels show the evolution of the energy density of each component. In the upper panels, the black solid lines show the temperature evolution, and the gray dashed lines indicate the analytic equilibrium temperature when the system evolves while conserving its total energy (see text). In the lower panels, the blue solid line represents the total energy (gas plus radiation), the orange and green dashed lines denote the gas and radiation energy densities, respectively, and the red dotted line indicates the radiation energy density corresponding to the Planck distribution, $aT^{4}$.
@@
As discussed in Section III.1, the use of the RSLA can lead to violations of energy conservation during the energy exchange between gas and radiation, making it difficult to accurately follow the evolution of optically thick, radiation-dominated systems. To address this issue, in Section IV.1, we have introduced a non-RSLA radiation component, $E^{\mathrm{N}}_{\mathrm{rad}}$, which is not subject to the RSLA and becomes dominant in optically thick regions such that $\tau_{\mathrm{cell}}\gtrsim 2/(3f_{c})$. To verify the effectiveness of this approach, we evolve a single computational cell over time under local emission and absorption processes, without including hydrodynamics or RT. This test demonstrates that our new scheme conserves the total energy and accurately reproduces the thermal equilibrium state expected under optically thick conditions.
@@
VI.1.2 Radiative transfer test with the diffusion approximation
Figure 2: Radiative transfer test with the diffusion approximation. The blue solid line represents the spherically averaged radial profile in the steady-state from our simulation. The orange dashed line denotes the analytic solution. The gray dotted line indicates the computational cell size (6.25 au), which corresponds to the finite volume over which radiation is injected. The simulation is performed in a cubic box of size 200 au with a uniform grid of $32^{3}$.
@@ VI.1.2 Radiative transfer test with the diffusion approximation
When radiation from a central source propagates through a uniform medium, the steady-state solution can be obtained analytically. In the optically thick limit, radiation propagates with diffusion as $F=-c\nabla E_{\mathrm{rad}}/(3\rho\sigma^{\nu}_{0})$. On the other hand, assuming a steady state, the flux can be expressed with the luminosity of the central radiation source $L$ as $F=L/(4\pi r^{2})$. As mentioned in Rosdahl and Teyssier (2015), combining the two expressions for the flux yields the radiation energy density of
@@
VI.1.3 Advecting radiation pulse test
Figure 3: Advecting radiation pulse test. The top and middle panels show the distributions of density and temperature. The blue solid and orange dashed lines represent the static and moving cases, respectively, while the gray dotted lines denote the initial conditions. The bottom panel presents the relative errors between the static and moving cases. The green and red lines indicate the relative errors in density and temperature, respectively. The simulation is performed in a cubic box of size 1024 cm with a uniform grid of $256^{3}$. All quantities are evaluated at $t=4.8\times 10^{-4}$ s.
@@ VI.1.3 Advecting radiation pulse test
where $T_{0}=10^{7}$ K, $T_{1}=2\times 10^{7}$ K, $\rho_{0}=1.2~\mathrm{g~cm}^{-3}$, $w=24$ cm, and $\mu=2.23$. The resulting initial profiles are shown as gray dotted lines in the top and middle panels of Figure 3. As the system evolves, the central radiation peak gradually broadens due to diffusion. We perform this test under two conditions: one in which the gas remains static, and another in which the gas moves uniformly at a constant velocity of $10^{5}~\mathrm{cm~s^{-1}}$. If the scheme accurately captures radiation advection in the dynamic diffusion limit, both cases should exhibit identical evolution in the comoving frame (Krumholz et al., 2007a). We set the computational domain to be a cubic box with a side length of $1024$ cm, resolved by $256^{3}$ uniform cells. Periodic boundary conditions are applied in all directions. We adopt a constant absorption opacity of $\kappa^{\nu}_{0}=10^{4}~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and neglect scattering. Under these conditions, the optical depth per cell is $\tau_{\mathrm{cell}}\gtrsim 4\times 10^{3}$, indicating that the radiation energy is dominated by the non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$.
@@
VI.1.4 Radiation pressure tube test
Figure 4: Radiation pressure tube test. In the top panel, the blue and orange lines show the density $\rho$ and temperature $T$ profiles along the $x$-axis after ten sound-crossing times, while the dashed and dotted lines represent the analytic hydrostatic equilibrium solutions for density and temperature, respectively. In the bottom panel, we show the relative errors of the numerical results presented in the top panel with respect to the analytic solutions. The blue and orange lines represent the relative errors in density and temperature, respectively. The simulation is performed in a cubic box of size 128 cm with a uniform grid of $128^{3}$.
Here, we demonstrate that our scheme accurately captures radiation pressure in the optically thick regime, where the non-RSLA component $E^{\mathrm{N}}_{\mathrm{rad}}$ dominates. To this end, we examine whether our scheme can maintain a hydrostatic structure supported by both gas and radiation pressure. First, to construct the initial condition for 3D simulations, we obtain the gas profile in steady-state hydrostatic equilibrium under optically thick conditions, where radiation is transported by diffusion, by solving the following 1D equations, which describe the so-called radiation pressure tube problem (Krumholz et al., 2007a):
@@ VI.1.4 Radiation pressure tube test
VI.2 Radiative transfer test with steep optical depth gradient
Figure 5: Radiative transfer test with steep optical depth gradient. We show spherically averaged radial profiles of outward luminosity in the steady state. The blue line represents the result obtained by accounting for the cell optical depth $\tau_{\mathrm{cell}}$ of adjacent cells, while the orange line shows the result obtained using only the local value of $\tau_{\mathrm{cell}}$. The horizontal dashed line marks the expected luminosity, $4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$. The vertical dotted line indicates the clump radius, $R_{\mathrm{c}}$. The simulation is performed in a cubic box of size $200$ au with a uniform grid of $128^{3}$.
As mentioned in Section IV.2, if radiation is divided into the RSLA and non-RSLA components solely based on the local value of $\tau_{\mathrm{cell}}$, it cannot escape from optically thick cells even when the adjacent cell is optically thin. To address this issue, we distribute the radiation components by also considering the $\tau_{\mathrm{cell}}$ values of neighboring cells, as detailed in Section V.3 and V.4. In this section, we verify that this method allows radiation to escape from optically thick regions into adjacent optically thin regions. To isolate the radiative behavior, we disable hydrodynamics and neglect scattering for simplicity.
We set up a computational box with a side length of $200~\mathrm{au}$, resolved with $128^{3}$ uniform cells. At the center of the box, we place a gas clump with radius $R_{\mathrm{c}}=30~\mathrm{au}$, density $\rho_{\mathrm{c}}=10^{-2}~\mathrm{g}~\mathrm{cm}^{-3}$, and temperature $T_{\mathrm{c}}=10^{5}~\mathrm{K}$. Outside the clump, we place a gas with a density of $\rho=10^{-20}~\mathrm{g}~\mathrm{cm}^{-3}$ and an initial temperature of $T=100~\mathrm{K}$. The gas temperature within the clump is kept fixed throughout the calculation. We set a constant absorption opacity of $\kappa^{\nu}_{0}=10^{2}~\mathrm{cm^{2}~g^{-1}}$. As the outer boundary condition for radiation, we adopt Equation (98) with $\alpha=-2$, corresponding to an optically thin radiation field whose intensity decreases as $r^{-2}$. Due to this opacity value, the optical depth per cell becomes extremely large inside the clump, reaching $\tau_{\mathrm{cell}}=5\times 10^{11}$, while the region outside the clump remains optically thin. We then follow the radiation emitted by the clump. In the ideal case, the clump radiates from its surface with a luminosity of $L=4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$.
@@ VI.2 Radiative transfer test with steep optical depth gradient
VI.3 Radiation spectrum test
In our scheme, the radiation spectrum is reconstructed from the photon mean energy, as described in Section IV.3. In this section, we validate the ability of our method to accurately track the evolution of the radiation spectrum. We set up a computational box with a side length of $320~\mathrm{au}$, resolved with $256^{3}$ uniform cells. Then, as illustrated by the dashed line in Figure 6, we place two clumps along the $x$-axis at positions $x=80$ and 80-80 au, with temperatures of $T_{\mathrm{c}}=6\times 10^{4}$ and $3\times 10^{4}$ K, and radii of $R_{\mathrm{c}}=10$ and $40$ au, respectively. Both clumps have identical luminosities, given by $L=4\pi R_{\mathrm{c}}^{2}\sigma_{\mathrm{SB}}T_{\mathrm{c}}^{4}$. The gas density inside each clump is set to $10^{-2}~\mathrm{g~cm}^{-3}$, while the ambient medium has a density of $10^{-16}~\mathrm{g~cm}^{-3}$ and an initial temperature of $100$ K. To isolate radiation effects, we turn off hydrodynamics and fix the gas temperature within the clumps. The opacity is set such that $\kappa^{\nu}_{0}=100~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and $\sigma^{\nu}_{0}=0~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ inside the clumps, and $\kappa^{\nu}_{0}=0~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ and $\sigma^{\nu}_{0}=100~\mathrm{cm}^{2}~\mathrm{g}^{-1}$ outside. This setup ensures that radiation is emitted only within the clumps, while outside it is scattered without absorption or emission. As the outer boundary condition for radiation, we adopt the same condition as in Section VI.1.2, which is appropriate for an optically thick radiation field. Starting from an initial condition with no radiation field, we evolve the system until it reaches a steady state.